OpenStudy (anonymous):

Absolute max/min for f(x)=ln(x+2)+1/x 1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

f'(x) = 1/x+2 - 1/x^2 = 0 1/x+2 = 1/x^2 x+2 = x^2 x^2 -x -2 = 0 (x-2)(x+1) = 0 x = 2, -1 -1 outside domain so disregard look at f' value for x less than 2 to determine concavity f'(1) = 1/3 - 1 = -2/3 slope neg so concave up gives min min f(x) is x=2

OpenStudy (anonymous):

wow thank you

OpenStudy (dumbcow):

your welcome actual answer is f(2), just plug in 2 for x

OpenStudy (anonymous):

Okay thanks agian